Griggs Company produces a single product with a current selling price of $170. Variable costs are $130 per unit, and fixed costs per month average $6,240. Management is considering increasing the selling price to a proposed $190 per unit. Assume that the variable cost per unit of the product and monthly fixed expenses will not change as a result of the proposed increase in selling price.

Hint: Treat each situation (current and proposed price) as separate potential scenarios when evaluating each question.

At the current selling price of $170 per unit, closest to what dollar volume of sales per month is required for Griggs to break-even? (Round the intermediate percentage to one decimal place, and round the answer up to the next whole unit.)

Multiple Choice

  • $20,800

  • $8,299

  • $26,554

  • $6,178
Answer #1

Contribution margin ratio = Contribution margin/Sales

= (170-130)/170 = 23.5%

Breakeven sales = Fixed cost/Contribution margin ratio

= 6240/23.5%

= 26,554

Option C is the answer